# Spectral Flow for the Riemann zeros

### v1

@@ II The model and it’s Bethe Ansatz equation
In the infinite interval for $E>2\pi e$ the above dispersion relation can be inverted as
@@
where $W$ is the principal branch of the Lambert $W$ function. Since $E(p)$ must be greater than $2\pi e$, this dispersion relation has an intrinsic energy gap. For large $p$, $E(p)\approx p/\log p$, which would be relativistic if it weren’t for the $1/\log p$. This is not a standard dispersion relation, however many uncommon dispersion relations arise in condensed matter physics.
@@
As for $p(E)$ in eq. (2), $\vartheta(E)$ is monotonic and positive in the infinite interval $(E_{*},\infty)$, where $E_{*}\simeq 17.8456$, and in this interval it is invertible and therefore asymptotically eq.(3) is valid. Thus (11) should be viewed as a large $t$ asymptotic version of the exact equation (17). However for reasons already mentioned, it is quite accurate for relatively low $t$, and (11) provides excellent results for even the first few zeros. For instance from (11) with $\delta=10^{-6}$ we obtain
@@
Since the S-matrix for the LM model, which is based on the Euler product formula, is unitary, the underlying hamiltonian is such that all eigen-energies $E_{n}(\sigma)$ are real in the thermodynamic limit with $\sigma>\tfrac{1}{2}$.
@@ VII Conclusions
We proposed that the quantized energies $\{E_{n}(\sigma)\}$ for the LM model are real due to the fact that its S-matrix is unitary based on the Euler product (Proposition 1.) Without assuming the RH, we studied the spectral flow, namely how the $E_{n}(\sigma)$ depend on $\sigma$ and derived a simple condition for all eigenvalues to be real, stated as Propositions 2 and 3 above. We showed how these Propositions are equivalent to the validity of the Riemann Hypothesis. We extended this reasoning to the two infinite classes of functions for the Generalized and Grand Riemann Hypotheses. We also presented spectral flow arguments for a counterexample where the RH is violated, and attributed this to the lack of an Euler product which implies a non-unitary S-matrix. This work provides ample evidence that for the RH to be true, one needs both the functional equation and the Euler product.
Based on this article, the validity of the RH comes down Proposition 1, namely that the hamiltonian for the LM model and its generalizations has real eigenvalues $E_{n}(\sigma)$. We argued that this is valid due to the unitarity of the S-matrix and the reality and invertibility of the dispersion relation. Defining the hamiltonian for the “easy” smooth part, namely the dispersion relation (2) which leads to the simple Lambert approximation (3), is not so obvious, and is left as an open question. A priori, we cannot think of any obstructions to establishing Proposition 1 for the ‘free”, i.e. non-interacting part, so long as it is invertible, which we have shown for large enough $t\gtrsim 10$. We point out that the recent work Negro seems relevant to this problem.

### v2

@@ II The model and it’s Bethe Ansatz equation
In the infinite interval for $E>E_{\rm gap}\equiv 2\pi e\approx 17.08$, the above dispersion relation can be inverted as
@@
where $W$ is the principal branch of the Lambert $W$ function. Since $E(p)$ must be greater than $2\pi e$, this dispersion relation has an intrinsic energy gap $E_{\rm gap}$. The gap is obvious from the small $p$ expansion,
$\dfrac{E}{2\pi e}=1+\dfrac{p}{2\pi e}-\dfrac{p^{2}}{8\pi^{2}e^{2}}+\dfrac{p^{3}}{12\pi^{3}e^{3}}+{\cal O}(p^{4}).$ (4)
This shows that above the gap $E(p)$ can be interpreted as corrections to a relativistic dispersion. For large $p$, $E(p)\approx p/\log p$, which would be relativistic if it weren’t for the $1/\log p$. This is not a standard dispersion relation, however many uncommon dispersion relations arise in condensed matter physics. For our purposes, (4) implies that the hamiltonian is hermitian since $p$ is a hermitian operator: in position space it is the usual $p=-i\hbar\partial_{x}$.
@@
As for $p(E)$ in eq. (2), $\vartheta(E)$ is monotonic and positive in the infinite interval $(E_{*},\infty)$, where $E_{*}\simeq 17.8456$ is a minor deformation of the gap $E_{\rm gap}=2\pi e$, and in this interval it is invertible and therefore asymptotically eq.(3) is valid. Thus (12) should be viewed as a large $t$ asymptotic version of the exact equation (18). However for reasons already mentioned, it is quite accurate for relatively low $t$, and (12) provides excellent results for even the first few zeros. For instance from (12) with $\delta=10^{-6}$ we obtain
@@
Since the S-matrix for the LM model, which is based on the Euler product formula, is unitary, and the non-interacting part of the hamiltonian is hermitian based on (4), all eigen-energies $E_{n}(\sigma)$ are real in the thermodynamic limit with $\sigma>\tfrac{1}{2}$.
@@ VII Conclusions
We proposed that the quantized energies $\{E_{n}(\sigma)\}$ for the LM model are real due to the fact that its S-matrix is unitary based on the Euler product and the hamiltonian for the free part, i.e. the dispersion relation, is hermitian (Proposition 1.) Without assuming the RH, we studied the spectral flow, namely how the $E_{n}(\sigma)$ depend on $\sigma$ and derived a simple condition for all eigenvalues to be real, stated as Propositions 2 and 3 above. We showed how these Propositions are equivalent to the validity of the Riemann Hypothesis. We extended this reasoning to the two infinite classes of functions for the Generalized and Grand Riemann Hypotheses. We also presented spectral flow arguments for a counterexample where the RH is violated, and attributed this to the lack of an Euler product which implies a non-unitary S-matrix. This work provides ample evidence that for the RH to be true, one needs both the functional equation and the Euler product.
Based on this article, the validity of the RH comes down Proposition 1, namely that the hamiltonian for the LM model and its generalizations has real eigenvalues $E_{n}(\sigma)$. We argued that this is valid due to the unitarity of the S-matrix and the reality and invertibility of the dispersion relation. In fact, it is clear from (4) that the free hamiltonian for our dispersion relation is a hermitian operator. We thus cannot see any obstructions to 1 at present. We point out that the recent work Negro seems relevant to the study of the free non-interacting part of the hamiltonian if the dispersion relation is based on the exact Riemann-Siegel $\vartheta$ function.